# ROADWork: A Dataset and Benchmark for Learning to Recognize, Observe, Analyze and Drive Through Work Zones

### v2

@@ Abstract
Perceiving and autonomously navigating through work zones is a challenging and underexplored problem. Open datasets for this long-tailed scenario are scarce. We propose the ROADWork dataset to learn to recognize, observe, analyze, and drive through work zones. State-of-the-art foundation models fail when applied to work zones. Fine-tuning models on our dataset significantly improves perception and navigation in work zones. With ROADWork , we discover new work zone images with higher precision (+32.5%) at a much higher rate (12.8$\times$) around the world. Open-vocabulary methods fail too, whereas fine-tuned detectors improve performance (+32.2 AP). Vision-Language Models (VLMs) struggle to describe work zones, but fine-tuning substantially improves performance (+36.7 SPICE).
Beyond fine-tuning, we show the value of simple techniques. Video label propagation provides additional gains (+2.6 AP) for instance segmentation. While reading work zone signs, composing a detector and text spotter via crop-scaling improves performance (+14.2% 1-NED). Composing work zone detections to provide context further reduces hallucinations (+3.9 SPICE) in VLMs. We predict navigational goals and compute drivable paths from work zone videos. Incorporating road work semantics ensures 53.6% goals have angular error (AE) <0.5°absent0.5°<\$\$< 0.5  ° (+9.9 %) and 75.3% pathways have $AE<$$$ 0.5  ° (+8.1 %).
@@ 1 Introduction
Navigating safely through work zones is a significant challenge for both autonomous and human-driven vehicles. In recent years, several reports suggest that self-driving cars are stumped by the simplest work zones [20, 65, 117, 72, 69, 48, 116, 66, 22]. In fact, even humans find driving through work zones challenging [91, 71, 99]. Since 2010, more than 700 people have died every year in the United States in accidents involving work zones [73, 93]. Although road work zones are relatively rare, learning to navigate them is essential for full vehicular autonomy.
@@
BDD100k [107] C 100K Images None
NuScenes [5] C 1000 Scenes (40K Frames) 5 Object Categories Boston and Singapore
Waymo [89] C 1150 Scenes None 3 US Cities
Mapillary [74] C 25K Images 2 Object Categories Around The World
Cityscapes [17] C 5000 Images 1 Object Category 27 Cities in Germany
WildDash [108] LT 1800 Images None Around The World
WildDash 2 [109] LT 5000 Images None Around The World
@@
CODA [53] LT 1500 Scenes (5937 Images) 6 Object Categories and few cities in China
SegmentMeIfYouCan [6] AN 327 Images (Eval) None Unknown
BDD-Anomaly [34] AN 810 Images (Eval) None Data from BDD100K
@@
Table 1: Comparing ROADWork With Other Driving Datasets/Benchmarks. Compared to common (C) datasets [107, 5, 74, 17], ROADWork has a wider variety of annotations for a rare driving scenario while being similar in size. Compared to long-tailed (LT) and anomalous (AN) datasets [108, 109, 53, 6, 34], ROADWork is larger with rich pixel, object and scene-level annotations across a broader range of tasks. Others focus on one or two aspects of self-driving but ROADWork holistically covers work zone perception and navigation.
Despite the research potential, few studies focus on this domain. Work zones fall into the long-tail of scene distributions for vehicular autonomy. Mining data for such long-tail situations is expensive, difficult, and labor-intensive [56]. Even large data sources contain few informative and diverse long-tailed samples. Thus, such data and methods are often trade secrets [103, 32]. For perspective, in this work, we semi-automatically filtered only 5078 keyframes (less than 0.01%) from nearly 45 million frames in one data source.
Both common [107, 74, 17] and long-tailed datasets [53, 108, 109] contain few work zones. Among common datasets, our analysis estimates that BDD100K [107] and Mapillary [74] collectively contain fewer than 1000 work zone images. Although long-tailed datasets such as CODA [53] contain bounding boxes for conventional work zone objects such as cones and work vehicles, they do not cover all work zone objects and do not provide any instance segmentations, fine-grained object attributes, scene descriptions, or drivable paths. Others such as WildDash [108, 109] contain images with heavy weather, lighting and geographic variations but largely annotate common Cityscapes [17] categories. Previous works contain little research on work zones, focusing mainly on individual issues and edge cases [68, 30, 29, 104, 76, 26, 25, 45, 86].
Because work zones are neglected and under-represented in existing datasets, even after training with massive amounts of data and compute, foundation models such as open vocabulary detectors [115, 102, 61], promptable segmentation models [46, 79], scene text models [81] and vision language models [52, 59, 50], do not perform well in this scenario. The key to improving self-driving in work zones lies in obtaining work zone driving data - even simple baselines such as Mask R-CNN [33] outperform foundation models [115, 102] when trained on a well-curated dataset.
@@
What kind of data is needed for driving in work zones? No two work zones are truly alike, with unique arrangements of barriers, cones, and signs. Navigating these work zones requires context-specific understanding of atypical object configurations that standard driving datasets fail to capture. We propose that understanding and navigating work zones mirrors human cognition, where seeing differs from observing, and knowing differs from understanding [88]. We address: Recognizing work zones and their constituent objects, Observing fine-grained details (e.g. sign text) that provide navigation instructions, Analyzing global scene context to capture spatial relationships, and Driving through work zones via pathway prediction.
@@
  • Beyond fine-tuning, we show the value of simple techniques: (a) video label propagation improves detection (+2.8 $AP$), (b) crop-rescaling improves reading sign texts (+14.2% 1-NED), (c) composing work zone detections with VLMs reduces hallucinations in scene descriptions (+3.9 SPICE), and (d) incorporating work zone semantics improves pathway prediction (+9.9% in goals with angular error <0.5°absent0.5°<\$\$< 0.5  °). Fine-tuning foundation models on the ROADWork dataset does not hurt performance on common driving situations allowing label unification. The ROADWork dataset enables studying other long-tailed scenarios such as geographic domain adaptation.
@@ 2 ROADWork Dataset
ROADWork consists of 4375 videos, 9650 richly annotated images, and 129017 images with passable trajectories. The videos were used to propagate labels [79] to 11959 additional keyframes to train our models. Like other recent papers for scaling data [46], we use bootstrapping to collect long-tailed work zone data. Our first bootstrapping stage involved manually capturing 2338 work zone images by driving in Pittsburgh to train our initial models. In our second bootstrapping stage, we manually curated 5078 keyframes after automatically mining nearly 100K potential images from Michelin Mobility Intelligence (MMI) Open Dataset [41] containing 45 million frames. From the selected keyframes, we obtained 4375 thirty-second videos of driving through work zones. All data is captured using smartphones with known intrinsics mounted inside the vehicle. In total, we estimate that our data covers around 5000 work zones (see Appendix B) from the first two data curation stages. Most of the results presented in the paper correspond to training using data from the first two stages, unless otherwise stated. In our third bootstrapping stage, we expanded to in-the-wild data sources. We mined, discovered and annotated 969 work zone images from BDD100K [107] and Mapillary [74] for in-the-wild evaluation. We also semi-automatically discovered images in bad weather and night conditions. We discovered 465 images by driving in Pittsburgh and other rural US areas, along with 800 images from a commuter bus that followed a fixed route in Pittsburgh over the course of two years.
Figure 3: Generating Driving Trajectory. Using driving video frames (a-c) as input, we estimated camera poses using COLMAP [85] (d-f), where camera poses are depicted as viewing frustums shown in red. These poses were projected onto the estimated ground plane to form the driven trajectory, then back-projected onto the frames as the future driving trajectory (g).
@@ 3 Recognizing Work Zones
Detecting And Segmenting Work Zone Objects. Foundation models for open vocabulary scene understanding [115, 111, 105, 112] have shown impressive zero-shot generalization on many datasets. No closed vocabulary detector exists for recognizing work zones objects, leading us to consider these open vocabulary foundation models. Although earlier datasets considered few work zone objects such as cones [74, 5], they did not consider all types of work zone objects, such as “Arrow Board”.
@@
OpenSeeD (COCO+O365) [111] 2.9 5.8 2.5 1.3 2.8 4.0
Detic (LVIS+I21K) [115] 4.2 6.3 4.6 2.1 5.5 6.0
OMG-LLaVA (Many Datasets) [112] 0.2 0.3 0.2 0 0.1 0.8
@@
Mask R-CNN [33] 29.9 48.3 32.7 16.0 32.9 43.8
Mask DINO [51] 36.4 55.2 40.7 20.2 37.4 52.2
Mask DINO [51] w/ Video Prop. [79] 39.0 58.8 43.4 22.1 40.0 55.0
Table 2: Segmenting Work Zone Objects. We train instance segmentation [33, 51] on ROADWork using a coarse vocabulary. Open vocabulary methods [111, 115, 61] fail to recognize work zone objects. In contrast, supervised models perform significantly better (+32.2 $AP$), and label propagation through video [79] further improves performance (+2.6 $AP$).
We investigate whether popular open vocabulary instance segmentation methods, Detic [115] and OpenSeeD [111] generalize to our data. We also investigate whether OMG-LLaVA [112], a foundation model trained for pixel level tasks, can reason about and segment work zone objects. Table 2 shows our results. Although few of the object categories are common concepts, such as “Fence” and “Police Officer,” open vocabulary methods [115, 111] fail to detect and segment most objects, such as “Temporary Traffic Control (TTC) Message Board” and “Arrow Board”. This is irrespective of the large-scale training source. Note that all methods additionally leverage CLIP [77], which is trained on a 400 million image-text paired dataset, which likely under-represents work zones. For comparison, we train coarse vocabulary work zone object detectors by fine-tuning Mask R-CNN [33] and Mask DINO [51] on our dataset, producing significantly improved detection (+25.7 $AP$ and +32.2 $AP$, respectively). However, many work zone categories are rare, e.g. the ROADWork annotations itself has fewer than 1000 instances of “Arrow Board”. To increase the number of instances, we use the Segment Anything 2 (SAM2) [79] model to propagate segmentations through ROADWork videos – we prompt all ground truth masks as input for an annotated keyframe to SAM2. Then we propagate masks for the next 60 frames and uniformly add every 10th frame to our augmented dataset. We obtain 11959 more labeled frames using 1947 training videos from ROADWork , in addition to 5318 images in the training set. Training on this augmented dataset yields further improvements (+2.6 $AP$).
In summary, work zone objects are woefully underrepresented in existing foundation models and datasets, highlighting the crucial role of ROADWork dataset. We perform analysis of open-vocabulary detectors [61] and state-of-the-art object detectors in Appendix C.1, which follow similar trends. In Appendix C.1, we also show that these foundation models can be fine-tuned on our data with no or minimal loss in performance on common driving datasets.
$AP$ $AP50$ $AP75$ $AP$ $AP50$ $AP75$
Supervision SAM [46] SAM2 [79]
@@
Table 3: Are Manual Segmentations Still Necessary? We train instance segmentation models [33] with varying levels of weak supervision from the ROADWork dataset. We use SAM [46] and SAM2 [79] to generate pseudo-ground truth masks from ground truth boxes and points. Manual ground truth masks (last row) outperform SAM at tighter IoU thresholds (+5.4 $AP75$) and for rare categories such as “Arrow Board” (+26.9 $AP$). SAM2 pseudo-ground truth masks, while better than SAM psuedo-ground truth overall (+3.5 $AP$), perform worse at tighter thresholds (+6.3 $AP75$) when compared to manual ground-truth masks.
Are Manual Segmentations Still Necessary? Given foundation models for promptable segmentation [46, 79] have shown competitive zero-shot generalization on many datasets, are manual segmentations even needed anymore?
We use Segment Anything Model (SAM) [46] and Segment Anything Model 2 (SAM2) [79] to perform zero-shot instance segmentation. We use ground truth boxes as input to generate pseudo-ground truth masks for the ROADWork dataset. We also simulate $K=\{5,10\}$ positive and negative points from manual ground-truth masks as input prompts [13]. We then train detectors (in this case, Mask R-CNN [33]) using manual and pseudo-ground truth masks.
@@
Detic [115] I21K + LVIS 32 52.4 % 125 42.9 %
Mask R-CNN [33] COCO + ROADWork 411 84.9 % 558 77.0 %
Table 4: Characterizing and Discovering Work Zones. The ROADWork dataset improves discovery of new work zones compared to open-vocabulary foundation models. Using the same classification rule, we employ a coarse-vocabulary detector [33] trained on ROADWork dataset. We discovered 12.8$\times$ and 4.5$\times$ more work zone images than Detic [115], improving precision by +32.5% and +34.0% on BDD100K [107] and Mapillary [74].
Figure 5: Work Zones Discovered Around The World. Mapillary [74] dataset contains driving images from around the world. Despite the ROADWork training dataset images being restricted to the U.S., we discovered work zones captured in Europe and Asia.
@@
We discover work zones in existing self-driving datasets devising a simple classification rule: an image is a work zone iff at least three instances of work zone objects from two unique categories that occupy at least $10\%$ of the image area. We chose Mask R-CNN for discovery due to its good accuracy-latency trade-offs [54, 38] for processing millions of images. Compared to foundation model Detic [115], Mask R-CNN trained on ROADWork discovers work zone images with higher precision in BDD100K [107] and Mapillary [74] datasets. We found fewer than 1000 roadwork images, representing only $0.4\%$ and $2.3\%$ of these datasets respectively. Surprisingly, while the ROADWork dataset contains images only from US cities, the model discovered work zones worldwide (Figure 5). Furthermore, detectors trained on the ROADWork dataset perform significantly better than open-vocabulary methods on discovered images in a variety of other conditions along with BDD100K and Mapillary; see Appendix C.1.
Summary. The ROADWork dataset significantly helps to recognize and discover work zones. Further, In Appendix C.1, open-vocabulary detectors [61] fine-tuned on the ROADWork data still generalize to common driving scenes, while not degrading performance on Cityscapes [17] (+0.2 $AP$) compared to a pretrained model. ROADWork enables study of other interesting long-tailed recognition problems. In Appendix C.1.1, we show that there exist large domain gaps (+7.4 $AP$) between cities even with SOTA adaptation methods [43] (-13.8 $AP_{50}$ for “barricade” compared to upper-bound). We also show results for label unification. Detector performance degrades (-5.0 $AP$) when models are jointly trained on datasets with common and rare object labels. The state-of-the-art label-unification algorithm [102] marginally improves performance (+2.0 $AP$).

### v3

@@ Abstract
[…]
Beyond fine-tuning, we show the value of simple techniques. Video label propagation provides additional gains (+2.6 AP) for instance segmentation. While reading work zone signs, composing a detector and text spotter via crop-scaling improves performance (+14.2% 1-NED). Composing work zone detections to provide context further reduces hallucinations (+3.9 SPICE) in VLMs. We predict navigational goals and compute drivable paths from work zone videos. Incorporating road work semantics ensures 53.6% goals have angular error (AE) <0.5°<\$\$ (+9.9 %) and 75.3% pathways have AE<0.5°AE<\$\$ (+8.1 %).
@@ 1 Introduction
Navigating safely through work zones is a significant challenge for both autonomous and human-driven vehicles. In recent years, several reports suggest that self-driving cars are stumped by the simplest work zones [SelfDrivingChallenge2, news-wired2017construction, news-atlantic2023sf, news-pntv2023construction, news-cnn2021cones, news-nyt2023concrete, news-sfstandard2023hole, news-wired2022cruise, news-2024missionlocal]. In fact, even humans find driving through work zones challenging [Thapa2024, Morgan2010, Wang1996]. Since 2010, more than 700 people have died every year in the United States in accidents involving work zones [NSCInjuryFacts2023, NHTSAPeopleStats2021]. Although road work zones are relatively rare, learning to navigate them is essential for full vehicular autonomy.
@@
BDD100k [yu2020bdd100k] C 100K Images None
NuScenes [nuscenes] C 1000 Scenes (40K Frames) 5 Object Categories Boston and Singapore
Waymo [Waymodataset] C 1150 Scenes None 3 US Cities
Mapillary [neuhold2017mapillary] C 25K Images 2 Object Categories Around The World
Cityscapes [cordts2016cityscapes] C 5000 Images 1 Object Category 27 Cities in Germany
WildDash [zendel2018wilddash] LT 1800 Images None Around The World
WildDash 2 [zendel2022wilddash2] LT 5000 Images None Around The World
@@
CODA [li2022coda] LT 1500 Scenes (5937 Images) 6 Object Categories and few cities in China
SegmentMeIfYouCan [chan2021segmentmeifyoucan] AN 327 Images (Eval) None Unknown
BDD-Anomaly [hendrycks2019bddanomaly] AN 810 Images (Eval) None Data from BDD100K
@@
Table 1: Comparing ROADWork With Other Driving Datasets/Benchmarks. Compared to common (C) datasets [yu2020bdd100k, nuscenes, neuhold2017mapillary, cordts2016cityscapes], ROADWork has a wider variety of annotations for a rare driving scenario while being similar in size. Compared to long-tailed (LT) and anomalous (AN) datasets [zendel2018wilddash, zendel2022wilddash2, li2022coda, chan2021segmentmeifyoucan, hendrycks2019bddanomaly], ROADWork is larger with rich pixel, object and scene-level annotations across a broader range of tasks. Others focus on one or two aspects of self-driving but ROADWork holistically covers work zone perception and navigation.
Despite the research potential, few studies focus on this domain. Work zones fall into the long-tail of scene distributions for vehicular autonomy. Mining data for such long-tail situations is expensive, difficult, and labor-intensive [liang2024aide]. Even large data sources contain few informative and diverse long-tailed samples. Thus, such data and methods are often trade secrets [tesla-patent, cruise-long-tail]. For perspective, in this work, we semi-automatically filtered only 5078 keyframes (less than 0.01%) from nearly 45 million frames in one data source.
Both common [yu2020bdd100k, neuhold2017mapillary, cordts2016cityscapes] and long-tailed datasets [li2022coda, zendel2018wilddash, zendel2022wilddash2] contain few work zones. Among common datasets, our analysis estimates that BDD100K [yu2020bdd100k] and Mapillary [neuhold2017mapillary] collectively contain fewer than 1000 work zone images. Although long-tailed datasets such as CODA [li2022coda] contain bounding boxes for conventional work zone objects such as cones and work vehicles, they do not cover all work zone objects and do not provide any instance segmentations, fine-grained object attributes, scene descriptions, or drivable paths. Others such as WildDash [zendel2018wilddash, zendel2022wilddash2] contain images with heavy weather, lighting and geographic variations but largely annotate common Cityscapes [cordts2016cityscapes] categories. Previous works contain little research on work zones, focusing mainly on individual issues and edge cases [mathibela2013roadwork, gumpp2009recognition, graf2012probabilistic, wimmer2009automatic, pannen2020keep, ferguson2015construction, ferguson2015mapping, kim2024rosa, shi2021work].
Because work zones are neglected and under-represented in existing datasets, even after training with massive amounts of data and compute, foundation models such as open vocabulary detectors [zhou2022detecting, wang2023detecting, liu2024grounding], promptable segmentation models [kirillov2023segany, ravi2024sam2], scene text models [ronen2022glass] and vision language models [li2022blip, liu2023llava, li2024llavanext-strong], do not perform well in this scenario. The key to improving self-driving in work zones lies in obtaining work zone driving data - even simple baselines such as Mask R-CNN [he2017mask] outperform foundation models [zhou2022detecting, wang2023detecting] when trained on a well-curated dataset.
@@
What kind of data is needed for driving in work zones? No two work zones are truly alike, with unique arrangements of barriers, cones, and signs. Navigating these work zones requires context-specific understanding of atypical object configurations that standard driving datasets fail to capture. We propose that understanding and navigating work zones mirrors human cognition, where seeing differs from observing, and knowing differs from understanding [su2011glancing]. We address: Recognizing work zones and their constituent objects, Observing fine-grained details (e.g. sign text) that provide navigation instructions, Analyzing global scene context to capture spatial relationships, and Driving through work zones via pathway prediction.
@@
  • Beyond fine-tuning, we show the value of simple techniques: (a) video label propagation improves detection (+2.8 $AP$), (b) crop-rescaling improves reading sign texts (+14.2% 1-NED), (c) composing work zone detections with VLMs reduces hallucinations in scene descriptions (+3.9 SPICE), and (d) incorporating work zone semantics improves pathway prediction (+9.9% in goals with angular error <0.5°<\$\$). Fine-tuning foundation models on the ROADWork dataset does not hurt performance on common driving situations allowing label unification. The ROADWork dataset enables studying other long-tailed scenarios such as geographic domain adaptation.
@@ 2 ROADWork Dataset
ROADWork consists of 4375 videos, 9650 richly annotated images, and 129017 images with passable trajectories. The videos were used to propagate labels [ravi2024sam2] to 11959 additional keyframes to train our models. Like other recent papers for scaling data [kirillov2023segany], we use bootstrapping to collect long-tailed work zone data. Our first bootstrapping stage involved manually capturing 2338 work zone images by driving in Pittsburgh to train our initial models. In our second bootstrapping stage, we manually curated 5078 keyframes after automatically mining nearly 100K potential images from Michelin Mobility Intelligence (MMI) Open Dataset [roadbotics-open] containing 45 million frames. From the selected keyframes, we obtained 4375 thirty-second videos of driving through work zones. All data is captured using smartphones with known intrinsics mounted inside the vehicle. In total, we estimate that our data covers around 5000 work zones (see Appendix B) from the first two data curation stages. Most of the results presented in the paper correspond to training using data from the first two stages, unless otherwise stated. In our third bootstrapping stage, we expanded to in-the-wild data sources. We mined, discovered and annotated 969 work zone images from BDD100K [yu2020bdd100k] and Mapillary [neuhold2017mapillary] for in-the-wild evaluation. We also semi-automatically discovered images in bad weather and night conditions. We discovered 465 images by driving in Pittsburgh and other rural US areas, along with 800 images from a commuter bus that followed a fixed route in Pittsburgh over the course of two years.
Figure 3: Generating Driving Trajectory. Using driving video frames (a-c) as input, we estimated camera poses using COLMAP [schoenberger2016sfm] (d-f), where camera poses are depicted as viewing frustums shown in red. These poses were projected onto the estimated ground plane to form the driven trajectory, then back-projected onto the frames as the future driving trajectory (g).
@@ 3 Recognizing Work Zones
Detecting And Segmenting Work Zone Objects. Foundation models for open vocabulary scene understanding [zhou2022detecting, zhang2023simple, xu2023open, zhang2024omgllava] have shown impressive zero-shot generalization on many datasets. No closed vocabulary detector exists for recognizing work zones objects, leading us to consider these open vocabulary foundation models. Although earlier datasets considered few work zone objects such as cones [neuhold2017mapillary, nuscenes], they did not consider all types of work zone objects, such as “Arrow Board”.
@@
OpenSeeD (COCO+O365) [zhang2023simple] 2.9 5.8 2.5 1.3 2.8 4.0
Detic (LVIS+I21K) [zhou2022detecting] 4.2 6.3 4.6 2.1 5.5 6.0
OMG-LLaVA (Many Datasets) [zhang2024omgllava] 0.2 0.3 0.2 0 0.1 0.8
@@
Mask R-CNN [he2017mask] 29.9 48.3 32.7 16.0 32.9 43.8
Mask DINO [li2023maskdino] 36.4 55.2 40.7 20.2 37.4 52.2
Mask DINO [li2023maskdino] w/ Video Prop. [ravi2024sam2] 39.0 58.8 43.4 22.1 40.0 55.0
Table 2: Segmenting Work Zone Objects. We train instance segmentation [he2017mask, li2023maskdino] on ROADWork using a coarse vocabulary. Open vocabulary methods [zhang2023simple, zhou2022detecting, liu2024grounding] fail to recognize work zone objects. In contrast, supervised models perform significantly better (+32.2 APAP), and label propagation through video [ravi2024sam2] further improves performance (+2.6 $AP$).
We investigate whether popular open vocabulary instance segmentation methods, Detic [zhou2022detecting] and OpenSeeD [zhang2023simple] generalize to our data. We also investigate whether OMG-LLaVA [zhang2024omgllava], a foundation model trained for pixel level tasks, can reason about and segment work zone objects. Table 2 shows our results. Although few of the object categories are common concepts, such as “Fence” and “Police Officer,” open vocabulary methods [zhou2022detecting, zhang2023simple] fail to detect and segment most objects, such as “Temporary Traffic Control (TTC) Message Board” and “Arrow Board”. This is irrespective of the large-scale training source. Note that all methods additionally leverage CLIP [radford2021learning], which is trained on a 400 million image-text paired dataset, which likely under-represents work zones. For comparison, we train coarse vocabulary work zone object detectors by fine-tuning Mask R-CNN [he2017mask] and Mask DINO [li2023maskdino] on our dataset, producing significantly improved detection (+25.7 $AP$ and +32.2 $AP$, respectively). However, many work zone categories are rare, e.g. the ROADWork annotations itself has fewer than 1000 instances of “Arrow Board”. To increase the number of instances, we use the Segment Anything 2 (SAM2) [ravi2024sam2] model to propagate segmentations through ROADWork videos – we prompt all ground truth masks as input for an annotated keyframe to SAM2. Then we propagate masks for the next 60 frames and uniformly add every 10th frame to our augmented dataset. We obtain 11959 more labeled frames using 1947 training videos from ROADWork , in addition to 5318 images in the training set. Training on this augmented dataset yields further improvements (+2.6 $AP$).
In summary, work zone objects are woefully underrepresented in existing foundation models and datasets, highlighting the crucial role of ROADWork dataset. We perform analysis of open-vocabulary detectors [liu2024grounding] and state-of-the-art object detectors in Appendix C.1, which follow similar trends. In Appendix C.1, we also show that these foundation models can be fine-tuned on our data with no or minimal loss in performance on common driving datasets.
APAP $AP50$ $AP75$ APAP $AP50$ $AP75$
Supervision SAM [kirillov2023segany] SAM2 [ravi2024sam2]
@@
Table 3: Are Manual Segmentations Still Necessary? We train instance segmentation models [he2017mask] with varying levels of weak supervision from the ROADWork dataset. We use SAM [kirillov2023segany] and SAM2 [ravi2024sam2] to generate pseudo-ground truth masks from ground truth boxes and points. Manual ground truth masks (last row) outperform SAM at tighter IoU thresholds (+5.4 $AP75$) and for rare categories such as “Arrow Board” (+26.9 APAP). SAM2 pseudo-ground truth masks, while better than SAM psuedo-ground truth overall (+3.5 APAP), perform worse at tighter thresholds (+6.3 $AP75$) when compared to manual ground-truth masks.
Are Manual Segmentations Still Necessary? Given foundation models for promptable segmentation [kirillov2023segany, ravi2024sam2] have shown competitive zero-shot generalization on many datasets, are manual segmentations even needed anymore?
We use Segment Anything Model (SAM) [kirillov2023segany] and Segment Anything Model 2 (SAM2) [ravi2024sam2] to perform zero-shot instance segmentation. We use ground truth boxes as input to generate pseudo-ground truth masks for the ROADWork dataset. We also simulate $K=\{5,10\}$ positive and negative points from manual ground-truth masks as input prompts [cheng2022pointly]. We then train detectors (in this case, Mask R-CNN [he2017mask]) using manual and pseudo-ground truth masks.
@@
Detic [zhou2022detecting] I21K + LVIS 32 52.4 % 125 42.9 %
Mask R-CNN [he2017mask] COCO + ROADWork 411 84.9 % 558 77.0 %
[…]
Figure 5: Work Zones Discovered Around The World. Mapillary [neuhold2017mapillary] dataset contains driving images from around the world. Despite the ROADWork training dataset images being restricted to the U.S., we discovered work zones captured in Europe and Asia.
@@
We discover work zones in existing self-driving datasets devising a simple classification rule: an image is a work zone iff at least three instances of work zone objects from two unique categories that occupy at least $10\%$ of the image area. We chose Mask R-CNN for discovery due to its good accuracy-latency trade-offs [li2020towards, huang2017speed] for processing millions of images. Compared to foundation model Detic [zhou2022detecting], Mask R-CNN trained on ROADWork discovers work zone images with higher precision in BDD100K [yu2020bdd100k] and Mapillary [neuhold2017mapillary] datasets. We found fewer than 1000 roadwork images, representing only $0.4\%$ and $2.3\%$ of these datasets respectively. Surprisingly, while the ROADWork dataset contains images only from US cities, the model discovered work zones worldwide (Figure 5). Furthermore, detectors trained on the ROADWork dataset perform significantly better than open-vocabulary methods on discovered images in a variety of other conditions along with BDD100K and Mapillary; see Appendix C.1.
Summary. The ROADWork dataset significantly helps to recognize and discover work zones. Further, In Appendix C.1, open-vocabulary detectors [liu2024grounding] fine-tuned on the ROADWork data still generalize to common driving scenes, while not degrading performance on Cityscapes [cordts2016cityscapes] (+0.2 $AP$) compared to a pretrained model. ROADWork enables study of other interesting long-tailed recognition problems. In Appendix C.1.1, we show that there exist large domain gaps (+7.4 $AP$) between cities even with SOTA adaptation methods [kennerley20232pcnet] (-13.8 $AP_{50}$ for “barricade” compared to upper-bound). We also show results for label unification. Detector performance degrades (-5.0 $AP$) when models are jointly trained on datasets with common and rare object labels. The state-of-the-art label-unification algorithm [wang2023detecting] marginally improves performance (+2.0 $AP$).